Problem 1-35 Recording events in a horizontal statements model LO 1-3, 1-4, 1-5, 1-6, 1-7,1-8, 1-9, 1-10 The following information applies to the questions displayed below. Maben Company was started on January 1, 2018, and experienced the following events during its first year of operation: 1 Acquired $38,000 cash from the issue of common stock 2. Borrowed $32,000 cash from National Bank 3. Earned cash revenues of $56,000 for performing services 4. Paid cash expenses of $49,000 5. Paid a $1,800 cash dividend to the stockholders 6. Acquired an additional $28,000 cash from the issue of common stock 7. Paid $8,000 cash to reduce the principal balance of the bank note. 8. Paid $61,000 cash to purchase land. 9. Determined that the market value of the land is $85,000 Problem 1-35 Part f f. Determine the percentage of assets that were provided by investors, creditors, and earnings. (Round your answers to 2 decimal places.) Assets Investors Creditors Eamings
Answer #1

Considering the events given in the question, there are two assets available with the company i.e. cash and Land. To calculate cash balance, we need to prepare cash account.

Cash Account

Common stock 38,000 Expenses 49,000
Notes payable 32,000 Cash dividend 1,800
Service revenue 56,000 Note payable 8,000
Common stock 28,000 Land 61,000
Balance c/d 34,200
154,000 154,000

Hence, total assets = Cash + Land

= 34,200 + 61,000

= $95,200

Retained earnings = Service revenue - Cash expenses - Cash dividend

= 56,000 - 49,000 - 1,800

= $5,200

Assets of $95,200 were provided by investors, creditors and earnings as under:

Assets provided by investors i.e. common stockholders = $66,000

Hence, percentage of assets provided by investors = 66,000/95,200

= 69.33%

Assets provided by creditors i.e. notes payable = $24,000

Hence, percentage of assets provided by investors = 24,000/95,200

= 25.21%

Assets provided by earnings = $5,200

Hence, percentage of assets provided by earnings = 5,200/95,200

= 5.46%

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.